OpenStudy (anonymous):

Dan confirms the authenticity of an amethyst stone by a few tests. During one of the tests, he finds the angle of the refracted ray of light coming out of the amethyst into the air as 28°. If the refractive index of amethyst is 1.55, what is the incident angle? A) 152 B) 26 C) 47 D) 72

OpenStudy (anonymous):

C.47 by Snell's law\[\frac{ \sin \theta _{1} }{ \sin \theta _{2} }=\frac{ \eta _{1} }{ \eta _{2} }\] \[\sin \theta _{2}=\sin \theta _{1}\times \frac{ \eta _{1} }{ \eta _{2} }\] \[\sin \theta _{2}=\sin 28\times \frac{ 1.55 }{ 1 }\] if we solve this we get around 46.69 i.e. 47.